OpenStudy (anonymous):

Solve 2(4 – b) = 2 − b –(b − 6)

OpenStudy (anonymous):

0 = 0

OpenStudy (anonymous):

ok first distribute 2 and -b 8-2b=2-b-b+6 now take all the b terms to one side -2b+2b=8-8 0=0 it means you can use any value of b!!!!!

OpenStudy (anonymous):

I think she wrote the equation wrongly, check it again.

OpenStudy (anonymous):

@senoritaliseth can you check your Question? did you write it correctly?

OpenStudy (anonymous):

its ok guys, i appreciate the help.. :) the answer was b=0

OpenStudy (anonymous):

answer is b=anything